OpenStudy (anonymous):

A single slit has a width of 12.7 µm. Monochromatic light from a source passes through the slit and a fourth dark fringe (on one side of the central maximum) appears at an angle of 10.0° as measured from the CM. Calculate the wavelength of light used. What colour is this?

OpenStudy (anonymous):

\[d \sin \theta = \lambda\] d is the width of the slit, \[\theta\] is the angle of incidence at which the minimum intensity occurs, and \[\lambda\]is the wavelength of the light\ Light color corresponds to wavelenghth